TCCON > Sites > Indianapolis > Instrument History

Instrument History

August, 2012

  • 2012-08-06: TCCON container arrives to Ivy Tech College, Lawrence Campus (ahead of schedule). Fence complete.
  • 2012-08-07: TCCON gets repositioned onto solid concrete blocks. Site setup.
  • 2012-08-08: Site setup. Electrical work complete. A/C work complete.
  • 2012-08-09: Site setup.
  • 2012-08-10: Phone & internet online. Met station, Sun tracker & dome base installed.
  • 2012-08-11: Completed wiring of Zeno box (metereological station) and azimuth and elevation of the sun tracker. Installation, connection, and operation of UPS, Herc & Herc6 computers. Additionally, we setup the barometer, Wi-Fi inside the container. Software installation on new laptop.
  • 2012-08-12: In the IFS-125: installed detectors, beam splitters, H-Cl reference cell, reconnected carriage wire steel cable. Refilled chiller, checked for proper operation and leaks. Beam splitter damaged during cleaning.
  • 2012-08-13: Router configured and operational. Webcam visible at http://influx.no-ip.org  . Turned on IFS125 and got an error light and audible warning.
  • 2012-08-14: Removed shipping struts from retroreflector trolley. Now IFS is operating properly. Installed up OPUS 6.5 on virtual XP (x86) running on Windows 7 (64bit) laptop with aid of Baron V. using Webex.

  • 2012-08-15: Sun tracker working fine. Offset values adjusted.

  • 2012-08-16: Sun tracker continued to work correctly. Geolocation values fine-tuned. 855ft / 982hPa.

  • 2012-08-21:

    • Changed filename to reflect the fact that the optical filter was removed.
    • OPUS ran 0.012 cm-1 res. measurement without error! 
    • Found that HCl cell was touching aperture wheel and needed to be turned 180 deg
    • The modulation efficiency is calculated according to this calculation, assuming the zero is the same at MOPD as it is for ZPD. 
    •   LASA PK-PK Zero ME LASB PK-PK Zero ME
      ZPD 0.73 V
       -7.35V
       8.08V  -8.88V 0.72  7.15V
      1.74V
       5.41V  0.08V  0.62
      MOPD  0.38V
      -7.08V
       7.46  -8.88V 0.67  6.95V
      1.90V
       5.05V  0.09V  0.58
       
    • After adjusting the fixed cube corner:
        LASA PK-PK Zero ME LASB PK-PK Zero ME
      ZPD 0.76 V
       -7.34V
       8.1V  -8.89V 0.72   7.16V
       1.71V
       5.45V  0.01V  0.62
      MOPD  0.41V
      -7.04V
       7.45  -8.92V 0.66  6.95V
      1.89V
       5.05V  0.03V  0.58
       
    • After adjusting the laser entrance mirror:
        LASA PK-PK Zero ME LASB PK-PK Zero ME
      ZPD  0.71V
       -7.32V
       8.03V  -8.88V 0.72   7.18V
       1.71V
       5.47V  0.04V  0.62
      MOPD  0.47V
      -7.10V
       7.57V  -8.94V 0.67  7.03V
       1.86V
       5.17V  0.04V  0.59
      So no improvement after either adjustment.
       
    • Did one last offset adjust of the suntracker with the alignment aparatus in place (-44.85, 124.25).  Removed the aperture and replaced window.  Not happy about the position of the sun on the floor of the trailer; it is about 2 cm offset north from the previous markings.  Remounted tube and 45 deg turning mirror. (-45.4, 123.95), adjusted quadrant sensors and put into tracking mode.  It clouded up, so will spend more time on this tomorrow.
    • InGaAs = 3.085 V optimized; Si = 0.507 V optimized (1.3 mm aperture)
    • Discovered that HeNe lamp was set to Intensity rather than frequency mode.  Turned off the laser and changed this.   Attempted to run a final lamp run, but OPUS kept failing after last scan with, "Peak location out of range, please save peak position: 367 Peak location is out of range, please save peak position 367".  (Number has varies between 367 and 404).  Tried going out of OPUS, resetting the FTS, switching computers (i.e. OPUS software), checked that laser intensities as the scanner moved between the front and back short adjust, and remeasured DC voltage on InGaAs.  Could not find a problem.
  • 2012-08-22:
    • So, after (re-)leveling the base of the solar tracker and repeating the leveling with the mirror and leveled plate on the floor, we still find that the resulting circle of light cast on the floor is displaced from the circle you drew on the tape by about 1.5 cm northwest). We have watched the new circle of light for about 1.5 hr and it has drifted very little (~0.5 cm).  Remounted CaF2 window and replaced tube/45 deg mirror. Verified that tube was level.  The tracker was able to track in the Position Sun mode, so switched to Tracking mode.  Finally, centered the beam on the tube with the quadrant sensor.
    • Still trying to solve the "Peak location out of range" problem.  Switched to OPUS 7.0 on Caltech laptop, power cycled FTS and restarted OPUS, swapped beamsplitters - to no avail.
    • Finally reached Baron, who informed us that we needed to click on the Save Peak Position in the Check Signal tab of the Measurement Menu of OPUS.  This value changed apparently as a result of changing the mode of the laser from Intensity to Frequency.  OPUS runs fine now.  First measurement (if20120822lgee0x.003) today was obtained with the "spare" CaF2 beamsplitter.  Repeated with the original CaF2 beamsplitter (if20120822lgee0x.004).
    • Recorded final mod. eff.:
        LASA PK-PK Zero ME LASB PK-PK Zero ME
      ZPD  0.44V
       -7.34V
       7.78V  -8.94V 0.71   7.05V
       1.68V
       5.37V  0.02V  0.62
      MOPD  0.17V
      -7.13V
       7.30V  -8.94V 0.67  6.90V
       1.81V
       5.09V  -0.01V  0.58
    • Centered the solar beam on the exit field stop (AP2) with the adjustable flat mirror just before AP2. Optimized DC value of the InGaAs detector = -1.758 V; Si detector = -0.238 V (APT=1000).
    • After several confusing hours dealing with sporadic OPUS errors, FTS beeping, and OPUS measurements stopping before all scans completed, we stopped running OPUS 6.5 on Jim's virtual Windows XP machine and switched to OPUS 7.0 on Caltech's laptop (the Windows 7 version). The problems disappeared and we have made some progress on the coaligning of the interferometer and IR. We explred both horizontal and vertical adjustments and stopped finally to look at Linefit.  (Completed linefit, but need to transfer data to Mac to run Matlab later in the morning (it's 1:15 a.m. and I need to sleep!)
    • We plumbed in pump, pumped FTS down to 0.7hPa, reset the ST hardware, and started IFSloop.  
  • 2012-08-23
    • All systems came up with IFSloop this morning!!!
    • Continued to coalign and run Linefit on all the results:
  • 2012-08-24
    • Ran the ghost test with filter (70-3F-FA) using the 7kHz XPM file dictated by Messershmidt et al.  The GPR's are shown in the table below:
      Scanner Frequency Number of Scans Ghost Amplitude Parent Amplitude GPR Parent Wavenumber Ghost Wavenumber
      7.5 kHz 113  0.00001675  0.33433  5.01x10-5  5913.277  9882.033
           0.000023097  0.333845  6.92x10-5  5913.205  9882.376
           0.000008851  0.333502  2.65x10-5  5913.221  9882.471
           0.000001775  0.332702  5.34x10-6  5913.031  9882.143
    • 1/4 CW, 1/2 CCW, 1/4 CCW - done!
    • Transferred the data point table (.dpt) files from the final vented and evacuated measurements take with OPUS to ccycle machine at Caltech to run them through the Linux version of Linefit.  Differences attributed to differences in settings/parameters of the two lamp measurements and running Linefit on the MAC versus Linux machine.
    • To correct the pressure for latitude and altitude, we measured the temp and pressure with the thermometer/barometer and compared this to the Setra pressure measurement recorded by the Zeno and appearing on console 1 of IFSloop. (Conversion spreadsheet. Lat. = 39.861389 N, Alt. = 270 m)  Temp = 29.4 C, Barometer = 744.4 mm Hg (740.45 mm Hg @ 0 C local station pressure or 987.185 hPa), 984.9 hPa (Setra)
    • The scanner failed at 6:30 p.m. (local time), but the reason was not obvious.  The FTS seemed to be happy in idle, so I sent a sw Status ReInit command.  It started up fine.  We pumped down the instrument and prepared to leave at 7:30 p.m. (local time).  Going out the door, I noticed that the laser indicator was flashing "acquiring".  IFSloop still running, but close to dusk.
  • 2012-08-25
    • IFSloop was acquiring data and laser was stable at 10:25 a.m. (local time).
    • Temp = 25.5  C, Barometer =  746.1 mm Hg (742.62 mm Hg @ 0 C local station pressure or  990.1 hPa), 988.3 hPa (Setra)
    • Reset netcam password
    • 7:40 PM Temp = 24.9  C, Barometer =  745.0 mm Hg (741.60 mm Hg @ 0 C local station pressure or  988,7 hPa), 985.8 hPa (Setra)
    • IFS pressure 1.02 hPa at 7:31 PM - IFS looks to be holding vacuum well. HeNe laser has been stable.
  • 2012-08-26
    • 11:30 AM Temp = 26.4  C, Barometer =  746.8 mm Hg (743.21 mm Hg @ 0 C local station pressure or  990.9 hPa), 988.8 hPa (Setra)
    • adjusted Netcam pointing slightly and cleaned front window
    • 11:40 AM EDT IFS pressure of 0.84 hPa

September, 2012

  • 2012-09-03
    • Zeno failure occurred ~ 7:30 a.m.  The log shows some unexpected data strings from Zeno.  The input we are receiving in this state from the Zeno is definitely aberrant:
      16:41:18 [ERROR] Zeno: Expected string ',' at column 72
      16:41:18 [ERROR] Zeno: Input was: '\r\n4251,12/09/02,16:50:29,3.2,20,0.0,3,24.0,87,104.9,980.7,1,90,13.7,8000\x00,'
      16:41:21 [ERROR] Zeno: Expected string ',' at column 71
      16:41:21 [ERROR] Zeno: Input was: '\r\n4251,12/09/02,16:50:32,3.4,8,0.0,3,24.0,87,105.2,980.7,1,90,13.7,8000\x00'
      16:41:21 [ERROR] Zeno: Expected 4 digits at column 0
      16:41:21 [ERROR] Zeno: Input was: ','
      16:41:22 [ERROR] Zeno: Expected string ',' at column 73
      16:41:22 [ERROR] Zeno: Input was: '\r\n4251,12/09/02,16:50:34,3.4,333,0.0,3,24.0,87,105.2,980.7,1,90,13.6,8000\x00,'
      16:41:23 [ERROR] Zeno: Expected string ',' at column 72
      16:41:23 [ERROR] Zeno: Input was: '\r\n4251,12/09/02,16:50:35,3.4,10,0.0,3,24.0,87,
      105.3,980.8,1,90,13.7,8000\x00,'
      That '8000' is supposed to be the BIT data, which is reported in hex, and it is supposed to be followed by a comma. The output here in the log is escaped in order to identify non-printable characters received, and in this case we are receiving a NUL byte before the comma. I don't know if there is anything in the configuration of the Zeno that might cause that behavior, or if it is just a bug in the Zeno.

      Of course we have never really used the BIT data, so we could modify the driver to accept the record if everything looks good up to BIT and report 0 for the BIT data. It is also worth running the problem by the Zeno people to see if they can psych it out. 
  • 2012-09-04
    • There was no camera image this morning and the netcam was rebooted.
  • 2012-09-06
    • Found fiber cables for serial link between Zeno and Hercules dangling down from the ceiling. Suspected that this could be the cause of  the communications error. Remounted them to ceiling. Errors continue sporadically.
  • 2012-09-07
    • 11:40 AM EDT Temp = 23.5 C, Barometer = 740.5 mm Hg (737.29 mm Hg @ 0 C local station pressure or 982.96 hPa), 980.8 hPa (Setra)
    • There were Zeno errors throughout the night and morning. Did a sw Status Shutdown Quickly, then a qtalk -m /dev/ser1 to look at the Zeno. Several observations: Time is off by a little under 10 minutes. This I think is indicated by the Tdrift of -555 on the IFSloop Zeno column. The wind speed and direction would drop to 0.0 and 0.0 then back to regular values. This may be the correct way for the Zeno to respond. I never saw an obviuos error looking with qtalk at the Zeno output using the telebyte fiber link. However, when I went back to IFSloop, the Zeno communication would fail after some time.
    • Did a hard reboot of the Zeno by unplugging from the wall and the battery. When it rebooted, it gave a message about not being able to sample all the sensors in the alloted time. It suggested going from 1 second to 2 seconds or more. Also the BIT data went from something like 800000 to 00000000. Went back to IFSloop, and the Zeno communication error would reappear after some time.
    • Replaced the Telebyte optical/serial link with a null modem cable, did another hard reboot of the Zeno, and the stream to qtalk looked fine. Restarted IFSloop and no Zeno errors after several hours! The instrument briefly had enough sunlight for the dome to open, sun tracker to find the sun, and IFS125 to take data.
  • 2012-09-08
    • 10:20 AM EDT Temp = 22.5 C, Barometer = 739.8 mm Hg (736.59 mm Hg @ 0 C local station pressure or 982.03 hPa), 980.1 hPa (Setra)
    • Installed a copper clad steel grounding rod (5/8" x 8') and connected it to the container with a copper cable. Only about 40 mV AC between container and grounding rod before hookup, and only 0.11 ma AC flows through the connection. This is mostly to help protect TCCON in case of a lightning strike. (n.b. there was a lot of lightning last night!)
    • 06:00 PM EDT Temp = 24.9 C, Barometer = 739.9 mm Hg (736.52 mm Hg @ 0 C local station pressure or 981.94 hPa), 980.1 hPa (Setra)
    • TCCON seems to be working fine at COB.
    • Power failure about 7:30 PM! It was restored on or before midnight. Dome had closed for usual sunset conditions.
  • 2012-09-09
    • Eaton UPS had been completely run down, depowering the IFS125 + HeNe, solar tracker, router, netcam, and monitor. Herc6 was still running. Phone and internet were down as well.
    • Recovery: switched off IFS125, then restarted Eaton UPS. Waited until HeNe had stabilized, then turned on IFS125 and waited until it looked functional.
    • Next morning phone was back, but not internet. Herc6 showed IFSloop still running OK. Shutdown IFSloop, then did a hard reboot of sun tracker, then a QNX-browser procedure to get sun tracker to initialize and talk to Herc6. Checked Zeno using qtalk - it seemed to being working OK. Restarted IFSloop. At this point IFS125, sun tracker, and Zeno were operating correctly. Had to open dome using the hand control. After that the dome operated correctly. TCCOn functional except for internet.
    • About 5:20 PM EDT internet was restored. Herc sent out backlog of e-mails.
    • Dome closed and went home correctly at sunset.
    • The HeNe went from stable to acquiring at least twice during this day. It seemed to be related to closing the door. Thermal issue?
    • 16:10 eastern, temperature = 23.1 C, barometer = 742.9 mm Hg, Setra = 984.1 hPa
  • 2012-09-10
    • Laser was stable upon arrival at 9:30 AM EDT and all morning.
    • Dome may need to be trained to a slightly more positive angle, as shadow comes close to edge of mirror before correction. I don't want to change this riht now, though.
    • 9:30 AM EDT Temp = 20.7 C, Barometer = 748.5 mm Hg (745.59 mm Hg @ 0 C local station pressure or 994.04 hPa), 991.8 hPa (Setra) 
  • 2012-09-24
    • IFS125 failure. Indicated by failure to recover after software reset (e-mail 3:20 p.m. local time).  Corrected with a software reset. 

October, 2012

  • 2012-10-15
    • 11:25 eastern, temperature = 19.6 C, barometer = 740.3 mm Hg, Setra = 981.6hPa
  • 2012-10-22
    • Noted that camera image was blank; rebooted the camera software.
  • 2012-10-23
    • Noted that camera image was blank (9:45 a.m. PDT); rebooted the camera software.
  • 2012-10-28
    • Noted that camera image was blank; rebooted the camera software at 17:55 (local time).

December, 2012

  • 2012-12-08
    • 16:59 eastern, temperature = 16.9 C, barometer = 741.6 mm Hg, Setra = 983.6 hPa
  • 2012-12-09
    • 15:06 eastern, temperature = 17.3 C, barometer = 736.5 mm Hg, Setra = 976.4 hPa
  • 2012-12-10
    • 8:30 eastern, temperature = 12.9 C, barometer = 735.8 mm Hg, Setra = 976.2 hPa

Back to top

Tag page